OpenStudy (anonymous):

Solve x^2 + 10x + 7 = 0 by completing the square. (find the equation) (x+ 10)^2= 93?

OpenStudy (athe):

\[x^2+10x+7=0\rightarrow (x^2+2\cdot 5\cdot x +25) -25 +7=0\rightarrow (x+5)^2=18\]